OpenStudy (anonymous):

The length of each side of a square wooden box, in inches, is represented by the expression 5z4. The volume of the box, in cubic inches, is (5z4)3. Which simplified expression represents the volume of the box? 125z64 125z12 125z7 15z12

OpenStudy (anonymous):

\[(5z^4)^3=(5^3)(z^4)^3=125z^{4*3}=125z^{12}\]